OpenStudy (anonymous):

I REALLY NEED HELP Mike poured 15 cups of water into a rectangular container measuring 10 inches by 6 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. Part B: How many cubic inches of water was poured into the container? Part C: What was the height of the water in the container? Show your work. Measurement is given as length by width by height.

OpenStudy (anonymous):

You need to convert the centimeters into inches first. 9cm/2.54cm= total inches. Volume = l*w*h = 9*6*(9/2.54) inch^3 15cups*14.44 inch^3 = total volume of water in inches^3

OpenStudy (anonymous):

210 total volume of water in inches^3?

OpenStudy (anonymous):

firstly, make sure you use uniform units, for instance lets use cm 10*2.54 and 6*2.54 then find the volume i.e v=l*w*h v=[10*2.54]*[6.2.54]*9

OpenStudy (anonymous):

@Calvins The rest of the question is in imperial, that's not going to go well.

OpenStudy (anonymous):

im so confused.

OpenStudy (anonymous):

@nukka are right, but conversion units are given, we can always convert our last answer. what do yo think?

OpenStudy (anonymous):

part b: you simply multiply 15 by 14.44 =[15*14.44] cubic inch

OpenStudy (anonymous):

And you get 216.66?

OpenStudy (anonymous):

No one is helping me understand